OpenStudy (anonymous):

Solve Algebraically: 3/x+2 - 1/x = 1/5x

OpenStudy (aakashsudhakar):

(3/x+2) - (1/x) = (1/5x) 3 - (x+2/x) = (x+2/5x) 3x - (x+2) = (x(x+2)/5x) 5x(3x) - 5x(x+2) = x(x+2) 15x^2 - 5x^2 - 10x = x^2 + 2x 9x^2 - 12x = 0 3x(3x - 4) = 0 x = 0 x = 4/3

OpenStudy (tkhunny):

Is that \(\dfrac{3}{x} + 2\) as you have written it, or \(\dfrac{3}{x+2}\)? Is that \(\dfrac{1}{5}x\) as you have written it, or \(\dfrac{1}{5x}\)?